# This work was supported by project from DOE-CESER

### v1

@@ VI-C Circuit Switcher
Circuit switchers help to safely and reliably control the flow of power flow in the electric grid. Furthermore, they are also used to disrupt the fault induced current during faulty conditions. This isolates the faulty equipment, protects against short circuits and overloading conditions.
Impact of Climate Change: Circuit switcher’s function can be impacted by both extreme cold and heat. In case of extreme cold, the circuit switcher has to pass the ice breaking test which states that the rated ice-breaking ability are 9.5mm and 19mm [128]. The rated ice-breaking ability is the maximum thickness of ice deposited on a device that will not interfere with its successful opening or closing. Under certain circumstances, an ice storm can cause a deposit of ice of such thickness that overhead lines fail and operation of specific switching equipment is impaired [129]. Ice is produced naturally in two ways, they are 1) clear ice appears from rain falling through air with a temperature between 0°C (+32F) and –10°C (+14F)., and 2) Rime ice, characterized by a white appearance from the air entrapped during ice formation, appears from rain falling through air with a temperature below –10°C (+14 °F) or from condensation of atmospheric moisture on cold surfaces.
In case of extreme heat, the issue faced by the circuit switcher is derating. The rated continuous current is the maximum current that the switch can carry continously under normal operating conditions without exceeding the specified temperature limits. The specified temperature limits for non-enclosed are higher than enclosed switchers because the switchers that are enclosed has lower heat dissipation behavior [129].
@@ VI-D Voltage Regulator
Voltage fluctuations in distribution systems can occur due to various reasons like 1) daily or seasonal variations in power demand, 2) low demand to high demand, 3) load density along the feeder, 4) load balancing activities, and 5) unbalanced loads and feeders [132]
Impact of Climate Change:
@@
Voltage regulator’s standard design ratings like kVA, voltage, and current (air cooled system) are defined based on ambient temperature of 40°C. Furthermore, the average temperature for air cooled system cannot exceed 30°C over 24-hour period [133]. Also, the altitude consideration for rating design assume no more than 3,300 feet. This is because as altitude increases, the reduction is air density results in adverse effect on temperature rise and dielectric strength (Tab. III) [133]. Tab. IV shows the derating behavior of voltage regulator at different temperatures and altitudes.
@@
New York state reports regulators concerns either due to failure or maintenance issues. The most common failure mode is the failure of moveable or stationary contacts [134]. The survey results showed that 90% of regulators are less than 40 year old. The average age of spare regulator is 31 years. There are a total of 2980 operating units and 345 spares, out of which 1742 operating units and 199 spare units have associated age data. This is because manufacturers in earlier years did not provide manufacture dates on their nameplates. 95 regulator related events system wide over the last ten years (2000-2010). 56 out of 95 events were due to voltage regulator failure leading to 52,797 customer interruptions. This is a failure rate of 0.19% per year and is below the average yearly failure rate of 0.29% based on survey involving ten North American utilities [134]. The failure rate of a voltage regulator increases when the ambient temperature increases which also causes indirect impact on energy demand. This increase in energy demand shows a strong correlation with reactive power demand rise during summer peaks or heat waves resulting in malfunction [133]. Even though ambient temperature and altitude impacts the voltage regulator, often their failures are associated with combination of maintenance and climate related.
@@
The factors affecting voltage regulators overloading are the following [133]: 1) peak load; 2) tap position at peak load; 3) load prior to peak load; 4) tap position prior to peak load; 5) duration of peak load; 6) ambient temperature; 7) on-load tap-changer maximum through-current rating. All these factors are considered when defining the ratings of a voltage regulator. For example, Fig. 20 shows how the rated current changes for various combination of the above factor. The effect of the ambient temperature, peak load, duration and initial load are extrapolated from calculations found in the loading guides IEEE Std C57.91 and IEC 60076-7 [133]. IEEE Std C57.91 and IEC 60076-7 provide the best known general information for the loading of voltage regulators under various conditions [133].
@@
Figure 20: Impact on rated current of voltage regulator at a specific initial loading and ambient temperature for various percent regulation and peak load duration. It is observed that under different ambient temperatures, the rated current of regulator is impacted from [133].
Existing Methods to Measure the Impacts:
Based on the above discussion, a common method to measure the impacts of ambient temperature on the voltage regulator can be monitoring of ambient temperature in conjunction with the peak demand to ensure the regulator’s derated design values are respected. This ensures the regulator functions within its design capabilities without a breakdown.
Existing Methods to Mitigate the Impacts: To mitigate the impact of extreme heat on the operation of voltage regulator, the following measures can be taken ensure overloading of regulator: 1) The voltmeters and other instruments used with regulators should be as accurate as the regulator, 2) If the regulator can’t be verified in neutral (lossy operating condition generates heat [133]), report the situation to the system operator, 3) Follow switching procedures: When switching between substations, follow the procedure in the Substation Switching Procedures Section, 4) Using a slightly larger bandwidth can reduce the number of operations, and 5) while the above points directly or indirectly increases the temperature in the regulator, the derating behavior of the equipment due to ambient temperature also must be considered when following the above points.
@@ VI-E Frequency Changer
Traditionally, frequency changers are electromechanical machines comprised of motor-generator set. But with the advent of solid state electronics, electronic frequency changers also became a possibility. Electronic frequency changers consist of a rectifier stage (AC to DC) that changes the frequency to a desired value through inverter. For voltage conversion a transformer can be used and it also provides galvanic isolation between input and output AC circuits. The electronic frequency changer has similar point of failure as any electronics-based equipment due to extreme heat from previous subsection discussions. Here, we discuss about the electromechanical-based frequency changers.
Impact of Climate Change: The impact on frequency changers can be assessed directly by understanding the impact of ambient temperature on motor and generator. The primary challenge is the derating. Ambient temperature can decrease the efficiency of generator-transformer set (GT) which further impacts the performance of the static transformer (ST). The manufacturer design ratings assume the operation is maintained with ambient temperatures below 40°C. However, the alternator derates by 3% for every 5°C increase in ambient temperature [135].
Existing Methods to Measure the Impacts: One approach to measure the impacts of extreme temperatures on the frequency changer by monitoring the ambient temperature immediately surrounding the equipment. Then referring the observed ambient temperature in the derating curve to understand the derated maximum current rating [136].Additionally, it is important to note the altitude since the altitude variations can result unusual ambient temperature effects on the alternator similar to voltage regulators.
Existing Methods to Mitigate the Impacts: Generally, these issues are resolved by placing the frequency changer in a climate control environment/enclosure. This is especially useful when the climate conditions can change drastically at a specific location.
@@ VI-F Substation Grounding Resistance
Substation ground resistance is the resistance in ohms between substation neutral and earth resistance (zero-potential reference). The NEC 250 recommends the grounding of panel and equipment to maintain a resistance of 25-ohms or less. In substations the grounding should maintain a resistance of 1 to 5-ohms, depending on size and voltage levels [137].
Impact of Climate Change: The literature discusses different approaches to measure the ground resistance. [138] provides an equation to calculate the substation ground resistance when grid area, soil resistivity and total length of buried ground rods are known. Substation ground resistance is given by
@@
where $R_{g}$ is the substation ground resistance in $\Omega$, $\rho$ is the soil resistivity in $\Omega m$, $A$ is the area occupied by the ground grid in $m^{2}$, and $L_{T}$ is the total buried conductor length in $m$.
@@
where $h$ is the depth of the ground grid in $m$. The soil resistivity ($\rho$) for different types of earth like wet organic soil, moist, soil, dry soil, and bedrock are given by $10$, $100$, $1,000$, and $10,000$ $\Omega m$ respectively [140] It can be observed from equations (1) and (2) that the substation ground resistance is directly dependent on the soil resistivity. Furthermore, soil resistivity has negative temperature coefficient which means as temperature decreases the soil resistivity increases [141]. For example, Tab. V [141] shows that when the temperatures are above freezing point, the change in resistivity for temperature variation is smaller when compared to the ambient temperature below freezing point. Furthermore, equation (2) also shows as the substation ground resistance is directly proportional to the soil resistivity ($\rho$) and inversely proportional to the depth of ground grid ($h$). Therefore, in colder climate conditions it is important to construct the ground resistance mesh deep inside the ground.
@@
Once construction of substation ground resistance is complete, it is challenging to measure the zero volts between ground resistance and earth. This is because of the sphere of influence of the grid. For large substations, measurements must be taken outside the sphere of influence which can mean long measurement traverses and in most cases, it is not practical [140].These measurements can be taken by a current probe and ground resistance measurement device. For example, IEEE 81:2012 [142] requires the current probes to be used at least 5 times the maximum diameter of the grounding system. Another approach is by utilizing the temperature impact on soil resistivity combined with the equations (1) and (2), one can estimate the impact of extreme temperatures on the ground resistance.
@@
Existing Methods to Mitigate the Impacts: [143] proposed a method to reduce the ground resistance values in three steps: (1) drilling deep holes in the ground; (2) developing cracks in the soil by means of explosions in the holes; (3) filling the holes with low resistivity materials (LRM) under pressure. Most of the cracks around the vertical conductors will be filled with LRM and a complex network of low resistivity tree-like cracks linked to the substation grid is formed.
VI-G Lightening Arrestor
@@
Lightening arrestor is a device that is used protect the insulation and conductors from lightening strikes. These devices divert the current from lightening surge through the arrestor into the ground.
Impact of Climate Change: Regular operation condition for lightening arrestors are 1) ambient temperatures between -40°C and 40°C, 2) altitudes of no more than 1000 $m$, and 3) wind speeds should be less than or equal to 34 $m/sec$ [144, 145].
@@
Existing Methods to Measure the Impacts: There are various monitoring methods used to monitor the health of lightening arrestors. On-line measurement methods are preferred since they do not require a substation to be taken out of service to perform the evaluations. Furthermore, unlike most of the equipment, lightening arrestor’s performance and health can only be monitored by comparing with the baseline measurements. Some of the online methods are 1) infrared imaging, 2) monitoring resistive component of current, 3) third harmonic measurement, 4) absolute temperature and comparison, and 5) power loss at maximum continuous over voltage [146].Monitoring of arrestors was dominated by third harmonic measurement (also known as leakage current test) at the base of the arrestor. However, this method requires the arrestor to be isolated from the earth for accurate data collection [147].To account for this, infrared imaging method is employed for online monitoring of arrestors. However, estimation of arrestor health using infrared imaging is not easy because of the lack of supplier’s information on absolute temperature correlation with arrestor’s health. For example, [148]shows how infrared scanning can help identify failures in work in three arrestors when compared to the each other. According to [147], a temperature of 5°C between two partner arrestors warrants concern and anything greater than 10°C means a failure. There are various failure modes for lightening arrestors, among them moisture ingress is known to be a major factor [146].[146] presents some field test showcasing the impact of different factors such as 1) time of the day, 2) location of sensor on arrestor, 3) cardinal direction of sun, and 4) weather on the absolute surface temperature of arrestor. These tests are conducted to identify the correlation of absolute surface temperature as metric for health monitoring instead of comparing a group of arrestors for health monitoring that is currently used in the field.
@@
Existing Methods to Mitigate the Impacts: The current online monitoring methods that has been in use for decades such as third harmonic measurement require isolation from earth and reliable data collection techniques while the infrared imaging method does not suffer from these issue. It suffers from standard method to accurately correlate absolute temperature to the arrestor’s health. The arrestor’s failure can only be mitigated by early detection.
@@ VI-H Relays
IEEE Std C37.91-2021 [149], IEEE Guide for Protecting Power Transformers, reports temperature standards about various other kinds of relays, that are currently being used commercially in the industry, as shown below -
@@
  • Adaptive Overcurrent Relays: In this type of relays, the pickup current value is obtained using inverse-time characteristic curve. This curve shows that the pickup current increases with decrease in ambient temperatures and vice versa. However, an exact formulation to modify the pickup current as a continuous function of the changing ambient temperature has not been suggested in this standard. It was also reported in [149] that in some utilities an alternate practice is being followed in which the pickup current is increased by 25% in the winter in severe climate swing locations.
@@
  • Gas accumulator relay: This relay, also known as Buchholz relay, is typically only used in transformers with conservator tanks due to no gas space being available inside the transformer tank. It is reported that these relays undergo heating issues due to increase in ambient temperature.
@@
  • Temperature relay: This relay, also known as the replica relay, typically measures and applies the phase current in the transformer to its heater units placed inside the relay. It was recommended that the installed relay should not be ambiently compensated and should experience same ambient temperature as the transformer it is protecting.
@@
Impact of Climate Change: Some of the typical challenges faced by relays due to changes in ambient temperature as follows. The relay coil in electromechanical relays is made of copper wire and the resistance of the coil increases by 0.4% per °C. This means as the temperature increases, the available current in coil to operate the contacts in the relay is not sufficient which results in failure of relays to operate when required [150], and is an example of derating. The Impact on the temperature derating factor (which proportionally decreases the maximum ambient temperature conditions by the manufacturers) for varying levels of altitude has been reported in [151]. The contacts of relay at their full rated load may cause ambient temperature to rise by $10\degree C$. For example, if a relay first switches at $20\degree C$, within the short duration, the ambient temperature can rise to $30\degree C$ [150, 152]. Therefore, the relays are impacted directly and indirectly through changes in ambient temperature. Reference [153] also showed that the functional coordination of components is impacted by the layout causing overheating of the components. [154] shows that the combination of high ambient temperatures, limited ventilation, and solar radiation can cause relay failures.
@@
Electromechanical relays use coils which are typically located near the conductor and are impacted by the ambient temperatures. However, the electromechanical are predominantly replaced by digital relays which use CTs and PTs as sensors in which case, the failure modes of digital relays follow the failure modes of CTs and PTs. The digital relays are also housed or enclosed and not exposed to ambient environment. It is to be noted that impact of ambient temperature on sensing technologies cause sensitivity and loss of coordination.
@@
Existing Methods to Mitigate the Impacts: The mitigation strategy to avoid the impact of ambient temperature on sensors is to adjust the pickup current based on the derating behavior of the device. This will help ensure the relays are operated successfully as designed for the system. This also ensure coordination of protection system.
For digital relays, there are new methods to determine the protection settings accurately based on data analytics that uses historical event dataset to update and develop better protection strategies aimed at extreme weather readiness [155]. There also active research in the field of setting-less protection that uses dynamic state estimation to determine the health of the system that the replay is protecting [156]. This can be particularly helpful during the extreme weather operations where relay coordination may tend to fail, however, this method can theoretically identify an unhealthy downstream system.
Another solution is to use predictive based methods. The approaches to predict the hottest spot temperatures and taking some preventive actions has also been discussed in [149]. It was suggested to use the equations reported in Appendix D in [149], to make these predictions, provided, the load current and ambient temperature remains the same in the foreseeable future. Typically, these predictions can be assumed to hold true for the next 30 mins and would need to be recalculated automatically every few minutes to account for the changes in ambient temperature after that short time window. Based on the predictions computed using this approach, many corrective or monitoring actions can be taken such as an automatic activation of full cooling system based on the predicted hotspot temperature rather than the actual hotspot temperature, alarm being triggered, the overtemperature or the time to excessive loss of life could be displayed continuously.

### v2

@@ VI-C Circuit Switcher
Circuit switchers help to safely and reliably control the flow of power flow in the electric grid. They are also used to disrupt fault-induced current during faulty conditions. This isolates the faulty equipment, protecting against short circuits and overloading conditions.
Impact of Extreme Temperature Operations: Circuit switcher’s function can be impacted by both extreme cold and heat.
In case of extreme heat, the issue faced by the circuit switcher is derating. The rated continuous current is the maximum current that the switch can carry continuously under normal operating conditions without exceeding the specified temperature limits. The specified temperature limits for non-enclosed are higher than enclosed switchers because the switchers that are enclosed have lower heat dissipation behavior [127].
@@ VI-D Voltage Regulator
Voltage fluctuations in distribution systems can occur for various reasons, including 1) daily or seasonal variations in power demand, 2) low demand to high demand, 3) load density along the feeder, 4) load balancing activities, and 5) unbalanced loads and feeders [130]
@@
Impact of Extreme Temperature Operations: Voltage regulators’ standard design ratings like kVA, voltage, and current (air cooled system) are defined based on an ambient temperature of 40°C. Further, the average temperature for an air-cooled system cannot exceed 30°C over a 24-hour period [131]. Also, the altitude consideration for rating design assume no more than 3,300 feet. This is because as altitude increases, the reduction is air density results in an adverse effect on temperature rise and dielectric strength (Tab. III) [131]. Tab. IV shows the derating behavior of voltage regulator at different temperatures and altitudes.
@@
New York state reports regulator concerns either due to failure or maintenance issues. The most common failure mode is the failure of movable or stationary contacts [132]. The survey results showed that 90% of regulators are less than 40 years old. The average age of a spare regulator is 31 years. There are a total of 2,980 operating units and 345 spares, out of which 1,742 operating units and 199 spare units have associated age data. This is because manufacturers in earlier years did not provide manufacture dates on their nameplates. There were 95 regulator-related events system wide over the last ten years (2000–2010). 56 out of 95 events were due to voltage regulator failure, leading to 52,797 customer interruptions. This is a failure rate of 0.19% per year and is below the average yearly failure rate of 0.29% based on survey involving ten North American utilities [132]. The failure rate of a voltage regulator increases when the ambient temperature increases, which also causes indirect impact on energy demand. This increase in energy demand shows a strong correlation with reactive power demand rise during summer peaks or heat waves resulting in malfunction [131]. Even though ambient temperature and altitude impacts the voltage regulator, their failures are often associated with a combination of maintenance and climate related factors.
@@
The factors affecting voltage regulators overloading are the following [131]: 1) peak load; 2) tap position at peak load; 3) load prior to peak load; 4) tap position prior to peak load; 5) duration of peak load; 6) ambient temperature; 7) on-load tap-changer maximum through-current rating. All these factors are considered when defining the ratings of a voltage regulator. For example, Fig. 20 shows how the rated current changes for various combination of the above factors. The effects of the ambient temperature, peak load, duration and initial load are extrapolated from calculations found in the loading guides IEEE Std C57.91 and IEC 60076-7 [131]. IEEE Std C57.91 and IEC 60076-7 provide the best known general information for the loading of voltage regulators under various conditions [131].
@@
Figure 20: Impact on rated current of voltage regulator at a specific initial loading and ambient temperature for various percent regulation and peak load duration. It is observed that under different ambient temperatures, the rated current of regulator is affected from [131].
Existing Methods to Measure the Impacts: Monitoring ambient temperature in conjunction with peak demand can ensure that the regulators’ derated design values are respected, i.e., the regulator is operating within its design capabilities.
Existing Methods to Mitigate the Impacts: The following measures can mitigate the impacts of extreme heat on voltage regulators: 1) Voltmeters and other instruments used with regulators should be as accurate as the regulator. 2) If the regulator cannot be verified in neutral (lossy operating condition generates heat [131]), report the situation to the system operator. 3) When switching between substations, follow the procedure in the Substation Switching Procedures section. 4) Use a slightly larger bandwidth to reduce the number of operations. 5) Where the above points directly or indirectly increase the temperature in the regulator, consider the derating behavior of the equipment due to ambient temperature.
@@ VI-E Frequency Changer
Traditionally, frequency changers are electromechanical machines composed of motor-generator sets. But with the advent of solid-state electronics, electronic frequency changers also became a possibility. Electronic frequency changers consist of a rectifier stage (AC to DC) that changes the frequency to a desired value through an inverter. For voltage conversion, a transformer can be used, and it also provides galvanic isolation between input and output AC circuits. Electronic frequency changers have similar vulnerabilities to extreme heat to other electronics-based equipment discussed in previous subsections. Therefore, this section focuses on electromechanical frequency changers.
Impact of Extreme Temperature Operations: The impact on frequency changers can be assessed directly by understanding the impact of ambient temperature on the motor and generator. The primary challenge is derating. Ambient temperature can decrease the efficiency of generator-transformer set (GT), which further affects the performance of the static transformer (ST). Manufacturer design ratings assume that operation is maintained at ambient temperatures below 40°C. However, the alternator derates by 3% for every 5°C increase in ambient temperature [133].
Existing Methods to Measure the Impacts: One approach to measure the impacts of extreme temperatures on the frequency changer is by monitoring the ambient temperature immediately surrounding the equipment, then referring the observed ambient temperature in the derating curve to understand the derated maximum current rating [134].Additionally, it is important to note the altitude. As in voltage regulators, altitude variations can result unusual ambient temperature effects on alternators.
Existing Methods to Mitigate the Impacts: Generally, these issues are resolved by placing the frequency changer in a climate-controlled environment or enclosure. This is especially useful when ambient conditions can change drastically at a specific location.
@@ VI-F Substation Grounding Resistance
Substation ground resistance is the resistance in ohms between substation neutral and earth resistance (zero-potential reference). The NEC 250 recommends the grounding of panel and equipment to maintain a resistance of 25 ohms or less. In substations the grounding should maintain a resistance of 1 to 5 ohms, depending on size and voltage levels [135].
Impact of Extreme Temperature Operations: The literature discusses different approaches to measure the ground resistance. [136] provides an equation to calculate the substation ground resistance when grid area, soil resistivity, and total length of buried ground rods are known. Substation ground resistance is given by
@@
where $R_{g}$ is the substation ground resistance in $\Omega$, $\rho$ is the soil resistivity in $\Omega$m, $A$ is the area occupied by the ground grid in m2, and $L_{T}$ is the total buried conductor length in m.
@@
where $h$ is the depth of the ground grid in m. The soil resistivity ($\rho$) for different types of earth like wet organic soil, moist, soil, dry soil, and bedrock are given by $10$, $100$, $1,000$, and $10,000$ $\Omega$m respectively [138] It can be observed from equations (1) and (2) that the substation ground resistance is directly dependent on the soil resistivity. Furthermore, soil resistivity has a negative temperature coefficient, which means that as temperature decreases, soil resistivity increases [139]. For example, Tab. V [139] shows that when the temperatures are above freezing point, the change in resistivity for temperature variation is smaller when compared to the ambient temperature below the freezing point. Equation (2) also shows that the substation ground resistance is directly proportional to the soil resistivity ($\rho$) and inversely proportional to the depth of ground grid ($h$). Therefore, in colder climate conditions, it is important to construct the ground resistance mesh deep inside the ground.
@@
Once construction of substation ground resistance is complete, it is challenging to measure the zero volts between ground resistance and earth. This is because of the sphere of influence of the grid. For large substations, measurements must be taken outside the sphere of influence, which can mean long measurement traverses. In most cases, it is not practical [138].These measurements can be taken by a current probe and ground resistance measurement device. For example, IEEE 81:2012 [140] requires the current probes to be used at least 5 times the maximum diameter of the grounding system. Another approach is by using the temperature impact on soil resistivity combined with the equations (1) and (2) to estimate the effect of extreme temperatures on the ground resistance.
@@
Existing Methods to Mitigate the Impacts: Ref. [141] proposed a method to reduce the ground resistance values in three steps: (1) drilling deep holes in the ground; (2) developing cracks in the soil by means of explosions in the holes; (3) filling the holes with low resistivity materials (LRM) under pressure. Most of the cracks around the vertical conductors will be filled with LRM, and a complex network of low resistivity tree-like cracks linked to the substation grid is formed.
VI-G Lightning Arrestors and Surge Arrestors
@@
Lightning arrestors protect equipment against damaging voltages when lightning strikes. These devices divert the current from lightning surge through the arrestor into the ground.
Impact of Extreme Temperature Operations: Regular operating conditions (from standards that govern performance) for lightning arrestors are 1) ambient temperatures between $-40$°C and 40°C, 2) altitudes of no more than 1000 m, and 3) wind speeds less than or equal to 34 m/sec [142, 143]. However, these equipment can be deployed in mountainous regions above 1000 m as well, in which case their derating nature needs to be considered.
@@
Existing Methods to Measure the Impacts: Various monitoring methods are used to monitor the health of lightning arrestors. On-line measurement methods are preferred, because they do not require a substation to be taken out of service to perform the evaluations. Furthermore, unlike equipment, lightning arrestor performance and health can only be monitored by comparing with the baseline measurements. Some of the online methods are 1) infrared imaging, 2) monitoring resistive component of current, 3) third harmonic measurement, 4) absolute temperature and comparison, and 5) power loss at maximum continuous over voltage [144]. Monitoring of arrestors was dominated by third harmonic measurement (also known as leakage current test) at the base of the arrestor. However, this method requires the arrestor to be isolated from the earth for accurate data collection [145]. To account for this, infrared imaging method is employed for online monitoring of arrestors. However, estimation of arrestor health using infrared imaging is not easy because of the lack of information from suppliers on absolute temperature correlation with arrestors’ health. For example, [146]shows how infrared scanning can help identify failures in work in three arrestors when compared to the each other. According to [145], a temperature of 5°C between two partner arrestors warrants concern and anything greater than 10°C means a failure. There are various failure modes for lightning arrestors; in particular, moisture ingress is known to be a major factor [144].Ref. [144] presents some field tests showcasing the impact of different factors such as 1) time of day, 2) location of sensor on arrestor, 3) cardinal direction of sun, and 4) weather on the absolute surface temperature of arrestor. These tests are conducted to identify the correlation of absolute surface temperature as metric for health monitoring instead of comparing a group of arrestors for health monitoring that is currently used in the field.
@@
Existing Methods to Mitigate the Impacts: Arrestor failure can only be mitigated by early detection. Current online monitoring methods that have been in use for decades, such as third harmonic measurement, require isolation from earth and reliable data collection techniques, while the infrared imaging method does not suffer from these issue. However, it suffers from a lack of standard methods to accurately correlate absolute temperature to the arrestors’ health.
@@ VI-H Relays
IEEE Std C37.91-2021 [147], IEEE Guide for Protecting Power Transformers, reports temperature standards about various other kinds of relays that are currently being used commercially in the industry, as shown below:
@@
  • Adaptive overcurrent relays: In this type of relay, the pickup current value is obtained using an inverse-time characteristic curve. This curve shows that the pickup current increases with a decrease in ambient temperatures and vice versa. However, an exact formulation to modify the pickup current as a continuous function of the changing ambient temperature has not been suggested in this standard. It was also reported in [147] that in some utilities, an alternate practice is being followed in which the pickup current is increased by 25% in the winter in severe climate swing locations.
@@
  • Gas accumulator relay: This type of relay, also known as a Buchholz relay, is typically only used in transformers with conservator tanks due to no gas space being available inside the transformer tank. It is reported that these relays undergo heating issues due to increase in ambient temperature.
@@
  • Temperature relay: This type of relay, also known as the replica relay, typically measures and applies the phase current in the transformer to its heater units placed inside the relay. It was recommended that the installed relay should not be ambiently compensated and should experience same ambient temperature as the transformer it is protecting.
@@
Impact of Extreme Temperature Operations: Some of the typical challenges faced by relays due to changes in ambient temperature as follows. The relay coil in electromechanical relays is made of copper wire, and the resistance of the coil increases by 0.4% per degree C. This means that as the temperature increases, the available current in the coil to operate the contacts in the relay is not sufficient, which results in failure of relays to operate when required [148]. This diminished capacity is an example of derating. The impact on the temperature derating factor (which proportionally decreases the maximum ambient temperature conditions by the manufacturers) for varying levels of altitude has been reported in [149]. The contacts of the relay at their full rated load may cause ambient temperature to rise by 10°C. For example, if a relay first switches at 20°C, within the short duration, the ambient temperature can rise to 30°C [148, 150]. Therefore, the relays are affected directly and indirectly through changes in ambient temperature. Ref. [151] also showed that the functional coordination of components is affected by the layout, causing overheating of the components. Ref. [152] shows that the combination of high ambient temperatures, limited ventilation, and solar radiation can cause relay failures.
@@
Electromechanical relays use coils that are typically located near the conductor and are affected by the ambient temperatures. However, most electromechanical relays have been replaced by digital relays, which use CTs and PTs as sensors. The failure modes of digital relays follow the failure modes of CTs and PTs. Digital relays are also housed or enclosed and not exposed to the ambient environment. Note that ambient temperature can cause sensitivity and loss of coordination in sensing technologies.
@@
Existing Methods to Mitigate the Impacts: The mitigation strategy to avoid the impact of ambient temperature on sensors is to adjust the pickup current based on the derating behavior of the device. This will help ensure that the relays operate as designed. This also ensure coordination of protection system.
For digital relays, there are new methods to determine the protection settings accurately based on data analytics. These methods use historical event datasets to update and develop better protection strategies aimed at extreme weather readiness [153]. There is also active research in the field of settingless protection that uses dynamic state estimation to determine the health of the system that the replay is protecting [154]. This can be particularly helpful during the extreme weather operations where relay coordination may tend to fail; however, this method can theoretically identify an unhealthy downstream system.
Another solution is to use predictive methods. The approaches to predict the hottest spot temperatures and take preventive action have been discussed in [147]. It was suggested to use the equations reported in Appendix D in [147], to make these predictions, provided that the load current and ambient temperature remains the same in the foreseeable future. Typically, these predictions can be assumed to hold true for the next 30 minutes and would need to be recalculated automatically every few minutes to account for the changes in ambient temperature after that short time window. Based on the predictions computed using this approach, many corrective or monitoring actions can be taken, such as an automatic activation of the full cooling system based on the predicted hotspot temperature rather than the actual hotspot temperature, an alarm being triggered, or displaying the overtemperature or the time to excessive loss of life.